### v2

@@ Abstract
We design a large-language-model (LLM) agent that extracts causal feedback fuzzy cognitive maps (FCMs) from raw text. The causal learning or extraction process is agentic both because of the LLM’s semi-autonomy and because ultimately the FCM dynamical system’s equilibria drive the LLM agents to fetch and process causal text. The fetched text can in principle modify the adaptive FCM causal structure and so modify the source of its quasi-autonomy–its equilibrium limit cycles and fixed-point attractors. This bidirectional process endows the evolving FCM dynamical system with a degree of autonomy while still staying on its agentic leash. We show in particular that a sequence of three finely tuned system instructions guide an LLM agent as it systematically extracts key nouns and noun phrases from text, as it extracts FCM concept nodes from among those nouns and noun phrases, and then as it extracts or infers partial or fuzzy causal edges between those FCM nodes. We test this FCM generation on a recent essay about the promise of AI from the late diplomat and political theorist Henry Kissinger and his colleagues. This three-step process produced FCM dynamical systems that converged to the same equilibrium limit cycles as did the human-generated FCMs even though the human-generated FCM differed in the number of nodes and edges. A final FCM mixed generated FCMs from separate Gemini and ChatGPT LLM agents. The mixed FCM absorbed the equilibria of its dominant mixture component but also created new equilibria of its own to better approximate the underlying causal dynamical system.
@@ 1 The Agentic Leash: Growing Fuzzy Cognitive Map Dynamical Systems from Text
These causal FCM feedback dynamical systems form local fuzzy or partial causal rules from the sampled documents. This local causal structure in turn defines global equilibrium limit cycles that serve as scenario-like answers to causal what-if questions. They also define the very source of the FCM dynamical system’s agency – its evolving equilibrium limit cycles. This differs from ordinary feedforward agentic systems whose agency resides only in programmed commands. Mixing FCMs can give both richer learned causal knowledge bases and richer global equilibria. The extraction process is agentic [acharya2025agentic] or partially autonomous because the FCM’s evolving global equilibria command the LLM agents to fetch and process further text that then tends to change the commanding FCM equilibria. Related work uses an autoencoder-like mapping to convert FCMs to text and continue the reverberatory process [panda2025causal]. This bidirectional process keeps the FCM’s LLM agents on a type of flexible agentic leash.
@@
Figure 1 shows a feedback causal sub-network of the complete 15-node FCM in Figure 2. An LLM agent grew the complete FCM from a recent AI article titled “ChatGPT Heralds an Intellectual Revolution” by Henry Kissinger et al. in the Wall Street Journal [kissinger2023chatgpt]. The green highlight shows just one embedded feedback loop that traces the causal flow from Human-AI Interaction to Human Cognition and then back to itself: Human-AI Interaction $\rightarrow$ Human Cognition $\rightarrow$ Human-AI Interaction. Here $\rightarrow$” denotes a negative edge and $\rightarrow$” denotes a positive edge. Even this smaller 5-node sub-FCM encodes equilibrium limit cycles that can serve as answers to what-if questions. Figure 7 below shows this process for a simpler FCM.
These learned FCM knowledge graphs consist of local causal rules. The rules give an immediate local form of interpretability or explainable AI (XAI)[8466590, guimera2020bayesian, horst2019explaining, samek2019explainable, giamattei2025causal]. The embedded or hidden limit cycles in feedback FCMs also give a global form of XAI.
@@
We need an LLM that can systematically extract causal variables and their causal relationships from the text without human supervision. These causal rules from the text should then predict the dynamics of the system described by the text even if the text does not mention the dynamics. FCMs can model this causal information as a weighted directed cyclic graph and also give qualitative equilibrium predictions from the equilibria of the approximating causal FCM [kosko1986differential, kosko1986fuzzy, kosko1988hidden, osoba2017fuzzy, ziv2018potential, glykas2010fuzzy, papageorgiou2013fuzzy, stach2010divide, kosko1988hidden, taber2007quantization].
@@
Figure 1 shows the simplest version of the FCM agentic set-up. An LLM takes the WSJ article “ChatGPT Heralds an Intellectual Revolution” by Henry Kissinger et alia as text input and extracts causal variables and causal relationships among those variables from the text. It then builds an FCM out of those causal rules. Figure 1 shows only 5 of the 15 nodes from the extracted FCM. Figure 2 shows the complete FCM.
@@
2 Extracting FCMs from Text and Related Work
@@
This paper shows how to extract FCM nodes and edges from raw text by just using the language of the text. It does not rely on labels or annotations. The FCM nodes are usually nouns or noun phrases associated with some qualitative or quantitative measure. Verbs in the text describe FCM causal edges. Edge extraction is transparent and interpretable because it gives direct evidence for the edges by quoting the text.
Prior work by Maryam Berijanian et al. [berijanian2024soft] also extracted FCMs from text sources using LLMs. But they trained their LLM through supervision and annotated their text input with (“source”, “target”, “direction”) tuples. This may help the LLMs extract FCMs but it defeats the whole purpose of automated FCM extraction because it requires a human to read the text and annotate the nodes.
Extracting FCM nodes from text somewhat resembles Named Entity Recognition (NER) tasks for LLMs[wang2023gpt]. Here the LLMs identify key information from the text and classify them based on predefined categories. The LLMs can extract nouns and noun phrases from a text based on the properties of nouns in that language. But every noun or noun phrase in the text does not correspond to a node in the FCM. That approach would grow impossibly large and inaccurate FCMs. Finely-tuned system instructions can prompt the LLM to pick only the nouns that also share the properties of an FCM node.
@@ 3 FCM Basics
3.1 Causal Edge Matrix
@@
3.2 Discrete Time Evolution of FCMs
@@ 5 Simulations
OpenAI’s ChatGPT-4.1 LLM agent respectively produced a 24-node FCM and a 20-node FCM based on the same unguided and guided prompts. Figure 11 shows the edge matrices correspoonding to these FCMs. The 15-node Gemini-FCM and the 20-node ChatGPT-FCM mixed with equal mixing weights to give the final 24-node mixed FCM. Figure 14 shows this FCM-mixture. Figure 13 shows its edge matrix while Figure 15 shows its limit cycle.
@@ 6 Conclusions
The mixing structure of FCMs also allows a given large text to break into chunks and subchunks with corresponding mixed FCM components. This process can help with node and edge identification in large documents. Mixing such FCMs still produces an FCM and this should further help the growth and uses of very-large-scale causal knowledge networks. More advanced agentic FCM systems will lengthen the agentic leash and allow the systems to grow and use complex causal networks as it both shapes and obeys the evolving global equilibria of the total FCM dynamical system.

### v3

@@ Abstract
We design a large-language-model (LLM) agent system that extracts causal feedback fuzzy cognitive maps (FCMs) from raw text. The causal learning or extraction process is agentic both because of the LLM’s semi-autonomy and because ultimately the FCM dynamical system’s equilibria drive the LLM agents to fetch and process causal text. The fetched text can in principle modify the adaptive FCM causal structure and so modify the source of its quasi-autonomy—its equilibrium limit cycles and fixed-point attractors. This bidirectional process endows the evolving FCM dynamical system with a degree of autonomy while the system still stays on its agentic leash. We show in particular that a sequence of three system-instruction sets guide an LLM agent as it systematically extracts key nouns and noun phrases from text, as it extracts FCM concept nodes from among those nouns and noun phrases, and then as it extracts or infers partial or fuzzy causal edges between those FCM nodes. We test this FCM generation on a recent essay about the promise of AI from the late diplomat and political theorist Henry Kissinger and his colleagues. This three-step process produced FCM dynamical systems that converged to the same equilibrium limit cycles as did the human-generated FCMs even though the human-generated FCM differed in the number of nodes and edges. A final FCM mixed generated FCMs from separate Gemini and ChatGPT LLM agents. The mixed FCM absorbed the equilibria of its dominant mixture component but also created new equilibria of its own to better approximate the underlying causal dynamical system.
@@ 1 The Agentic Leash: Growing Fuzzy Cognitive Map Dynamical Systems from Text
These causal FCM feedback dynamical systems form local fuzzy or partial causal rules from the sampled documents. This local causal structure in turn defines global equilibrium limit cycles that serve as scenario-like answers to causal what-if questions. They also define the very source of the FCM dynamical system’s agency – its evolving equilibrium limit cycles. This differs from ordinary feedforward agentic systems whose agency resides only in programmed commands. Mixing FCMs can give both richer learned causal knowledge bases and richer global equilibria. The extraction process is agentic [1] or partially autonomous because the FCM’s evolving global equilibria command the LLM agents to fetch and process further text that then tends to change the commanding FCM equilibria. Related work uses an autoencoder-like mapping to convert FCMs to text and continue the reverberatory process [13]. This bidirectional process keeps the FCM’s LLM agents on a type of flexible agentic leash.
@@
Figure 1 shows a feedback causal sub-network of the complete 15-node FCM in Figure 2. An LLM agent grew the complete FCM from a recent AI article titled “ChatGPT Heralds an Intellectual Revolution” by Henry Kissinger et al. in the Wall Street Journal [8]. We note that the first published FCM in the original FCM paper [10] described key variables of Mideast peace based on the author’s reading of a 1982 syndicated newspaper opinion piece from Henry Kissinger. The green highlight shows just one embedded feedback loop that traces the causal flow from Human-AI Interaction to Human Cognition and then back to itself: Human-AI Interaction $\rightarrow$ Human Cognition $\rightarrow$ Human-AI Interaction. Here $\rightarrow$” denotes a negative edge and $\rightarrow$” denotes a positive edge. Even this smaller 5-node sub-FCM encodes equilibrium limit cycles that can serve as answers to what-if questions. Figure 7 below shows this process for a simpler FCM.
These learned FCM knowledge graphs consist of local causal rules. The rules give an immediate local form of interpretability or explainable AI (XAI)[2, 6, 7, 15, 4]. The embedded or hidden limit cycles in feedback FCMs also give a global form of XAI.
@@
We need an LLM that can systematically extract causal variables and their causal relationships from the text without human supervision. These causal rules from the text should then predict the dynamics of the system described by the text even if the text does not mention the dynamics. FCMs can model this causal information as a weighted directed cyclic graph and also give qualitative equilibrium predictions from the equilibria of the approximating causal FCM [9, 10, 11, 12, 20, 5, 14, 17, 11, 18].
@@
Figure 1 shows the simplest version of the FCM agentic set-up. An LLM takes the WSJ article “ChatGPT Heralds an Intellectual Revolution” by Henry Kissinger et al as text input and extracts causal variables and causal relationships among those variables from the text. It then builds an FCM out of those causal rules. Figure 1 shows only 5 of the 15 nodes from the extracted FCM. Figure 2 shows the complete FCM.
@@
2 Using LLM Agents to Extract FCMs from Text
@@
This paper shows how to use LLM agents to extract FCM nodes and edges from raw text by using just the language of the text. It does not rely on labels or annotations.
The technique starts with the analysis of nouns and verbs in the target text. The FCM nodes are usually nouns or noun phrases. The noun terms can describe physical or social properties or policy variables or anything else that can exert or absorb causal effects. Verbs in the text describe FCM causal edges. The process of edge extraction is both transparent and interpretable because it gives direct evidence for the edges by quoting from the text itself.
Our automated agentic approach differs from other efforts to extract FCMs from text that use LLMs in some way. Schuerkamp [16] used LLMs and human input to extract causal edges from text using LLMs. They used evolutionary algorithms to learn the edge weights while their list of nodes needed human input. Berijanian [3] also extracted FCMs from text sources using LLMs. Their method relied on some initial human annotation of text with (“source”, “target”, “direction”) tuples.
Extracting FCM nodes from text somewhat resembles Named Entity Recognition tasks for LLMs[19]. Here the LLMs identify key information from the text and classify them based on predefined categories. The LLMs can extract nouns and noun phrases from a text based on the properties of nouns in that language. But every noun or noun phrase in the text does not correspond to a node in the FCM. That approach would grow impossibly large and inaccurate FCMs. Finely tuned system instructions can prompt the LLM to pick only the nouns that also share the properties of an FCM node.
@@ 3 FCM Basics
3.1 Causal Edge Matrix $E$
@@
3.2 Discrete-Time Evolution of FCMs
@@ 5 Simulations
OpenAI’s ChatGPT-4.1 LLM agent respectively produced a 24-node FCM and a 20-node FCM based on the same unguided and guided prompts. Figure 11 shows the edge matrices corresponding to these FCMs. The 15-node Gemini-FCM and the 20-node ChatGPT-FCM mixed with equal mixing weights to give the final 24-node mixed FCM. Figure 14 shows this FCM-mixture. Figure 13 shows its edge matrix while Figure 15 shows its limit cycle.
@@ 6 Conclusions
The mixing structure of FCMs also allows a given large text to break into chunks and subchunks with corresponding mixed FCM components. This process can help with node and edge identification in large documents. Mixing such FCMs still produces an FCM and this should further help the growth and uses of very-large-scale causal knowledge networks. More advanced agentic FCM systems will lengthen the agentic leash. This will allow the knowledge system to grow and use complex causal networks as the system both shapes and obeys the evolving global equilibria of the total FCM dynamical system.